OpenStudy (anonymous):

How would you prepare 2.00 L of a .250 M NaOH solution from 1.00 M stock solution of NaOH?

OpenStudy (abhisar):

We are given the molarity of a more concentrated solution (1.0 M stock solution) and the volume and molarity of a more dilute one containing the same solute (2 L of 0.250 M solution). We are asked to calculate the volume of the concentrated solution needed to prepare the dilute solution. Our plan is to use Equation \( \tt M_1V_1=M_2V_2\) where \(\tt M_1\) is the molarity of the more concentrated solution. Because \( \tt M_1V_1 = M_2V_2\), we can write \(\tt \Large V_1= \huge \frac{M_2\times V_2}{M_1} \\ \Large V_1= \huge \frac{2\times 0.250 }{1} = ?\) We see that if we start with \(\tt V_1~mL\) of 1.0 M NaOH and dilute it to a total volume of 2 L, the desired 0.250 M solution will be obtained. The calculated volume seems reasonable because a small volume of concentrated solution is used to prepare a large volume of dilute solution (that is, \(\tt V_1 < V_2\) and \( \tt M_1 > M_2\).